OpenStudy (anonymous):

does tan x ( 1 + cos x) equal sin x + (sin x/cosx)

OpenStudy (lgbasallote):

i guess this has to be done by proving?

OpenStudy (anonymous):

yep.

OpenStudy (lgbasallote):

take the right hand side \[\sin x + \frac{1}{\cos x}\] get the LCD then add

OpenStudy (anonymous):

tan x ( 1 + cos x) = tan x + tanx cosx= sin x/cos x + sinx

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

another way... \[\sin x + \frac{\sin x}{\\cos x} = \frac{\sin x \cos x + \sin x}{\cos x} = \frac{sin ( 1 + \cos x)}{\cos x} = \tan x (1 + \cos x)\]